OpenStudy (anonymous):

A racing car is mov ing round a circular track at a constant speed of 60 m S·I. The radius •. of the track is500m. What is rhe ratio ofthe centripetal force to the weight of the car?

OpenStudy (anonymous):

0.02 0.73 1.2 71.0 one of these is the answer but idk which one

OpenStudy (anonymous):

.73 is the answer. F=mv^2/r and W=mg . SO taking ratio of F and W as mentioned in the problem,we get F:W=v^2/r*g. F : W=.73 .

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Centripetal force= mv^2/r and weight=mg ratio=mv^2/r/mg=60*60/500/10 ans. 72.5

OpenStudy (anonymous):

You are welcome.